OpenStudy (anonymous):

Solve. x2 = 10x – 20

OpenStudy (anonymous):

Did you mean this? \[x^2 = 10x - 20\] \[x^2 -10x + 20 = 0\] \[x^2 -10x + 25 = 5\] \[ (x-5)(x-5) = 5 \] \[ (x-5)^2 = 5 \] \[ (x-5) = sqrt{5} \] Get this: \[x = 5 \pm \sqrt{5}\]

OpenStudy (anonymous):

Yeah, thank you (: